Find the area of the given triangle with angle B = 20°, angle C = 115°, and c = 16.
Answer by Alan3354(69443)

Angle A = 45 degs
Find side a using Law of Sines:
a/sin(45) = 16/sin(115)
a = 16*sin(45)/sin(115)
a is the base
----------------
h = 16*sin(20)
Area = bh/2 = 128*sin(20)*sin(45)/sin(115)
Area =~ 34.156 sq units